Find all real solutions of the quadratic equation.
(x)square + 2 x - 27= 0

Answer by Nate(3500)

x^2 + 2x - 27 = 0
x^2 + 2x = 27
(x + 1)^2 = 28
x = -1 +- sqrt(28)